# A joint optimization approach to identifying sparse dynamics using least squares kernel collocation

### v1

@@ 1 Introduction
The identification of ordinary differential equations (ODEs) and dynamical systems is a fundamental problem in control [gevers_personal_2006, ljung2010perspectives, ljungTheory], data assimilation [smith_application_1962, Jazwinski1970], and more recently in scientific machine learning (ML) [Brunton2016, rackauckas2020universal, Raissi2018Hidden]. While algorithms such as Sparse Identification of Nonlinear Dynamics (SINDy) and its variants [Kaptanoglu2022] are widely used by practitioners, they often fail in scenarios where observations of the state of the system are scarce, indirect, and noisy. In such scenarios modifications to SINDy-type methods are required to enforce additional constraints on the recovered equations to make them consistent with the observational data. Put simply, traditional SINDy-type methods work in two steps: (1) the data is used to filter the state of the system and estimate the derivatives, and (2) the filtered state is used to learn the underlying dynamics. In the regime of scarce, noisy and incomplete data, step 1 is inaccurate, which can propagate to poor results in the subsequent step 2.
In this paper, we propose an all-at-once approach to filtering and equation learning based on collocation in a reproducing kernel Hilbert space (RKHS) which we term Joint SINDy (JSINDy), and shows that the issues above can be mitigated by performing both steps together. This joins a broader class of dynamics-informed methods that integrate the governing equations directly into the learning objective, either as hard constraints or as least-squares relaxations, which couples the problems of state estimation and model discovery. Representative examples include physics-informed and sparse-regression frameworks based on neural networks, splines, kernels, finite differences, and adjoint methods [Raissi2019-so, Chen2021, ijcai2021p283, jalalian2025dataefficientkernelmethodslearning, Fung2025-odr, Hokanson2023-sr, rackauckas2020universal].
@@ 1.1 Problem setup
with known measurement matrices $\boldsymbol{V}_{n}$ (taken to be the identity when the entire state is observed), observation times $\boldsymbol{t}=(t_{n})_{n=1}^{N}$, and measurement noise $\boldsymbol{\epsilon}_{n}$ which are assumed to be independent with zero mean and bounded variance. Given these observations, our goal is to estimate the nonlinear map $f:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ from an instance of the $\boldsymbol{y}_{n}$, i.e., to find $\widehat{f}\approx f$, and simultaneously infer the process $\boldsymbol{x}(t)$, i.e., find $\widehat{\boldsymbol{x}}(t)\approx\boldsymbol{x}(t)$. Additionally, we ask that our representation of $\widehat{f}$ is sparse in some basis or library of functions $\Phi=\left\{\phi_{1},...,\phi_{Q}\right\}$. That is, we assume
@@
where only a small number of $\widehat{\theta}_{m,q}$ are nonzero in order to have an interpretable symbolic form, following the seminal work [Brunton2016, wang2011, rudy2016datadriven, schaeffer2017learning]. As we do not know a priori which features $\phi_{q}$ are active, we parametrize $\widehat{f}$ using the full feature library and promote sparsity in the coefficients $\widehat{\boldsymbol{\theta}}:=(\widehat{\theta}_{m,q}),m=1,\cdots,d,q=1,\cdots Q$.
@@
    Exacerbating this issue is the fact that the measurements $\boldsymbol{y}_{n}$ are sparse, indirect, and noisy, which leads to further challenges in estimating $x$ and its derivatives.
@@
We overcome the above challenges by proposing a joint estimation approach where an approximation $\widehat{\boldsymbol{x}}(t)$ to the true state $\boldsymbol{x}(t)$ is obtained at the same time as the approximation $\widehat{f}(t)$ to the nonlinear dynamics. The key idea in our approach is to enforce self consistency of the $\widehat{\boldsymbol{x}}$ and $\widehat{f}$, i.e., $\widehat{\boldsymbol{x}}$ should approximately satisfy the ODE defined by $\widehat{f}$ while matching the observed data. Since the data is limited and the problem is severely ill-posed, we propose an optimization approach to solve the problem with appropriate regularization terms for both the $\boldsymbol{\theta}$ parameters and the approximate state $\widehat{\boldsymbol{x}}$, in particular, we use a sparse regularizer for the former and a reproducing kernel Hilbert space (RKHS) penalty for the latter.
@@ 1.2 Summary of contributions
In view of SINDy and other existing works on equation discovery [bayesiansindy, Brunton2016, Fung2025-odr, Chen2021, ijcai2021p283, Hokanson2023-sr], we summarize our contributions as follows:
@@
    We develop a RKHS regularized least-squares approach to learning general ODEs directly from sparse and partial observations. (see Section 2).
@@ 1.3 Literature Review
Learning continuous time dynamical systems from time series data has been a prominent problem in system identification [aastrom1971system, ljung2010perspectives, keesman2011system]. Seminal works using symbolic regression [bongard2007automated, schmidt2009distilling] through genetic programming sought to discover physical laws from observational data. The works [wang2011, Brunton2016, rudy2016datadriven, schaeffer2017learning] developed and popularized the use of a sparsity prior over a feature library to promote parsimony in the discovered model, drawing on ideas from compressive sensing[donoho2006compressed, candes2006robust, candes2008intro]. Bayesian approaches have also been considered, which allow for uncertainty quantification and motivate new strategies for model selection [zhang2018robust, niven2019bayesian, yang2020bayesian, bayesiansindy, north2022bayesian, north2023bayesian].
@@
Many of these works consider a two-step framework which first estimates time-derivatives, then performs regression to estimate the dynamics. These approaches are straightforward to implement and often succeed when measurements of the process are sufficiently dense but fail when the data is scarcely sampled and noisy. These approaches first solve a smoothing problem to estimate the state and its derivatives, using methods such as total variation regularization [Chartrand2011-cv], kernel regression [LONG2024134095], Savitzky–Golay filters, smoothing splines [Varah1982-ay], or Kalman smoothing [haas2024kalman]. These approaches inevitably trade bias for robustness: as shown in [vanBreugel2020NumericalDifferentiation], noise-resistant estimators lie on a Pareto frontier balancing error magnitude and error correlation due to implicitly assuming an inaccurate underlying dynamical model. An alternative to estimating derivatives, but that still lies within a two-step framework is to adopt a weak-form formulation [Messenger2021a], which avoids explicit differentiation by matching integrals against test functions. However, this still requires estimating integrals from noisy and scarcely sampled data, a problem that remains challenging despite being more well posed than differentiation. Consequently, any two-step approach to learning dynamics is inherently limited by the absence of a proper dynamical model in its initial estimation stage.
@@
To address limitations of two-step approaches, dynamics-informed approaches jointly estimate the dynamics and the state, allowing each to reinforce the other. These include Physics-Informed Neural Networks-Sparse Regression (PINN-SR) [Chen2021] and Physics-Informed Spline Learning (PiSL) [ijcai2021p283] which use least squares collocation applied to neural networks and splines respectively, while obtaining sparsity in the coefficients of the ODE by alternating between applying variants of gradient descent and sequentially thresholded ridge regression. [jalalian2025dataefficientkernelmethodslearning] takes a least squares collocation approach using RKHS methods, while also representing the unknown differential equation in an RKHS. The ODR-BINDy method in [Fung2025-odr] applies a least squares relaxation to a finite difference discretization, and handles feature selection by greedily maximizing a Bayesian evidence over models defined by feature sets. [Ribera2022-nz] also applies a least squares relaxation to a finite difference discretization, but optimizes using a variational annealing approach, and focuses on identifying dynamics with hidden variables. Notably, our objective function (8) is closely related to that of [Fung2025-odr, Ribera2022-nz, ijcai2021p283, Lahouel2024] and [Chen2021], though we apply different discretization, variable selection, and optimization strategies. [Hokanson2023-sr] applies a finite difference discretization but imposes hard ODE constraints by solving the KKT system associated to the discretized problem. They promote sparsity via $\ell_{p}$ regularization, and optimize using iteratively reweighted least squares.
@@
Aside from [Hokanson2023-sr], which considers decreasing sampling frequencies, and [Ribera2022-nz], which considers hidden variables, none of these works deal with very low sampling frequencies, partial observations through general linear measurements, and higher order dynamics in the context of ODEs. These issues are, however, considered heavily in [Chen2021] and [jalalian2025dataefficientkernelmethodslearning] in the context of PDEs, where partial information on the state (being infinite dimensional) and higher order derivatives are fundamental to the problem. [Hokanson2023-sr] crucially point out that decoupling the level of discretization in their finite difference stencil from the sampling rate of the data is required for mitigating bias in estimating the ODE when the sampling rate of the data is low, and demonstrate this in a restricted setting on the the Van der Pol system. We generally consider sampling rates over an order of magnitude lower than other works and flexibly handle partial observations and higher order ODEs.
@@ 1.5 Notation
2 JSINDy: Joint State Estimation and System Identification
@@ 2.1 The abstract formulation
Our approach centers around the joint computation of the estimated state $\widehat{\boldsymbol{x}}(t)$, henceforth called state estimation, as well as the right hand side $\widehat{f}(x)$, henceforth called system identification. We will require that the pair $\widehat{\boldsymbol{x}}$ and $\widehat{f}$ are self consistent, in the sense that $\widehat{\boldsymbol{x}}$ should (approximately) satisfy the ODE defined by $\widehat{f}$:
@@
To this end, we impose self-consistency using a least squares penalty method [tristan_analysis, tristan_pde, sasha_tristan_partialmin, vanLeeuwen2013Mitigating, xu2025efficientkernelbasedsolversnonlinear, Raissi2019-so], which is discretized via collocation in an RKHS. Imposing that $\widehat{f}$ has a sparse representation in the library $\Phi=\{\phi_{1},\ldots,\phi_{Q}\}$, we adopt a parametrized notation for how it operates on $\boldsymbol{x}$ and its derivatives:
@@
For instance, if we consider the first order constant coefficient linear ODE $\frac{d}{dt}\boldsymbol{x}(t)=\boldsymbol{A}\boldsymbol{x}$, then we $D[\boldsymbol{x}](t)=\boldsymbol{x}$, and we would be looking for $\boldsymbol{\theta}=\boldsymbol{A}$.
@@
    Because we handle sparsity through the feature selection operator $\mathcal{S}$ we do not need to apply regularization to $\boldsymbol{\theta}$ to promote sparsity. In principal, however, the Frobenius norm $\frac{\mu}{2}\|\boldsymbol{\theta}\|_{F}^{2}$ could be replaced by a more general sparsity promoting regularizer $R(\boldsymbol{\theta})$ which would supplant $\mathcal{S}$.
@@
then the pair $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ is a fixed point of (6). This is a strong condition, but is likely to hold whenever $\mathcal{S}$ is a reasonably good sparsifier that is consistent with the collocation loss $\mathcal{L}_{C}$. This is also a desirable property in itself; we want for our estimate fixed to the active set to be best estimate possible.
The ADO algorithms of [ijcai2021p283, Chen2021] are exactly of the form (6) and motivate it as a surrogate to $\ell_{0}$ minimization. Similarly, the greedy evidence maximization in [Fung2025-odr] is related in that it solves a sequence of fixed-support nonlinear least squares problems, though in each iteration, separate subproblems are solved for each candidate feature to remove. This has the benefit of minimizing a fixed objective function, and ensuring convergence due to the monotone feature removal process. This is discussed in detail algorithmically in Section 3.1 and Section 3.3.
@@
To represent the $d$-dimensional trajectory $\boldsymbol{x}:[0,T]\to\mathbb{R}^{d}$, we work in a vector-valued RKHS $\mathcal{X}$ of functions $0,T]\mapsto\mathbb{R}^{d}$ generated by the $d$-fold Cartesian product of an RKHS $\mathcal{H}_{k}$, associated to a symmetric positive definite kernel $k$. This is equivalent to modeling each coordinate $x_{i}(t)$ of the trajectory $\boldsymbol{x}(t)$ separately in the RKHS $\mathcal{H}_{k}$, or by taking $\mathcal{X}$ to be the vector-valued RKHS with matrix valued kernel $K(t,t^{\prime})=k(t,t^{\prime})I_{d}$. We will require that for derivative orders j=0,,pj=0,...,p, and times $t\in[0,T]$, the derivative evaluation map $\boldsymbol{x}\mapsto\frac{d^{j}}{dt^{j}}\boldsymbol{x}(t)$ is a bounded linear operator on $\mathcal{X}$ (see Section A.2). It suffices to take $k$ such that $\mathcal{H}_{k}$ is compactly embedded in $C^{p}([0,T])$, which is satisfied if $k\in C^{p,p}([0,T]\times[0,T])$ [ZhouDerivative].
We choose the kernel $k$ to be the sum of a high-order Matérn kernel [Rasmussen2006Gaussian] and a constant kernel to account for systems where the time-average is nonzero. More precisely, with $h_{\nu}$ denoting the $\nu$-order Matérn covariance function and $\ell$ as a length scale parameter, we use kernels of the form
@@
where $K_{\nu}$ is the modified Bessel function of the second kind, and $c_{0},c_{1}>0$. These kernels induce RKHSs which satsify the boundedness condition of the derivative evaluation maps whenever $\nu>p+\frac{1}{2}$. See Section A.2 for more details on kernels and smoothness.
As a preliminary step in our model, we choose the kernel hyperparameters $c_{0},c_{1},\ell$, and the variance of the observation noise $\sigma^{2}=\mathbb{V}\left[\epsilon\right]$ using maximum likelihood and treat the distribution of $\epsilon$ as Gaussian, as is common in Gaussian process regression; see for example [Rasmussen2006Gaussian]. From there, $\|\boldsymbol{x}\|_{\mathcal{X}}^{2}$ corresponds to the regularization induced by the RKHS norm associated to this fitted Gaussian process.

### v2

@@ 1 Introduction
The identification of ordinary differential equations (ODEs) and dynamical systems is a fundamental problem in control [31, 57, 56], data assimilation [84, 41], and more recently in scientific machine learning (ML) [12, 71, 73]. While algorithms such as Sparse Identification of Nonlinear Dynamics (SINDy) and its variants [45] are widely used by practitioners, they often fail in scenarios where observations of the state of the system are scarce, indirect, and noisy. In such scenarios modifications to SINDy-type methods are required to enforce additional constraints on the recovered equations to make them consistent with the observational data. Put simply, traditional SINDy-type methods work in two steps: (1) the data is used to filter the state of the system and estimate the derivatives, and (2) the filtered state is used to learn the underlying dynamics. In the regime of scarce, noisy and incomplete data, step 1 is inaccurate, which can propagate to poor results in the subsequent step 2.
In this paper, we propose an all-at-once approach to filtering and equation learning based on collocation in a reproducing kernel Hilbert space (RKHS) which we term Joint SINDy (JSINDy), and show that the issues above can be mitigated by performing both steps together. This joins a broader class of dynamics-informed methods that integrate the governing equations directly into the learning objective, either as hard constraints or as least-squares relaxations, which couples the problems of state estimation and model discovery. Representative examples include physics-informed and sparse-regression frameworks based on neural networks, splines, kernels, finite differences, and adjoint methods [72, 22, 87, 40, 28, 38, 71].
@@ 1.1 Problem setup
with known measurement matrices $\boldsymbol{V}_{n}$ (taken to be the identity when the entire state is observed), observation times $\boldsymbol{t}=(t_{n})_{n=1}^{N}$, and measurement noise $\boldsymbol{\epsilon}_{n}$ which are assumed to be independent with zero mean and bounded variance. Given these observations, our goal is to estimate the nonlinear map $f:\mathbb{R}^{dp}\mapsto\mathbb{R}^{d}$ from an instance of the $\boldsymbol{y}_{n}$, i.e., to find $\widehat{f}\approx f$, and simultaneously infer the process $\boldsymbol{x}(t)$, i.e., find $\widehat{\boldsymbol{x}}(t)\approx\boldsymbol{x}(t)$. Additionally, we ask that our representation of $\widehat{f}$ is sparse in some basis or library of functions $\Phi=\left\{\phi_{1},...,\phi_{Q}\right\}$. That is, we assume
@@
where only a small number of $\widehat{\theta}_{m,q}$ are nonzero in order to have an interpretable symbolic form, following the seminal work [12, 97, 78, 81]. As we do not know a priori which features $\phi_{q}$ are active, we parametrize $\widehat{f}$ using the full feature library and promote sparsity in the coefficients $\widehat{\boldsymbol{\theta}}:=(\widehat{\theta}_{m,q}),m=1,\cdots,d,q=1,\cdots,Q$.
@@
    Exacerbating this issue is the fact that the measurements $\boldsymbol{y}_{n}$ are sparse, indirect, and noisy, which leads to further challenges in estimating $\boldsymbol{x}$ and its derivatives.
@@
We overcome the above challenges by proposing a joint estimation approach where an approximation $\widehat{\boldsymbol{x}}(t)$ to the true state $\boldsymbol{x}(t)$ is obtained at the same time as the approximation $\widehat{f}$ to the nonlinear dynamics. The key idea in our approach is to enforce self consistency of the $\widehat{\boldsymbol{x}}$ and $\widehat{f}$, i.e., $\widehat{\boldsymbol{x}}$ should approximately satisfy the ODE defined by $\widehat{f}$ while matching the observed data. Since the data is limited and the problem is severely ill-posed, we propose an optimization approach to solve the problem with appropriate regularization terms for both the $\boldsymbol{\theta}$ parameters and the approximate state $\widehat{\boldsymbol{x}}$, in particular, we use a sparse regularizer for the former and a reproducing kernel Hilbert space (RKHS) penalty for the latter.
@@ 1.2 Summary of contributions
In view of SINDy and other existing works on equation discovery [37, 12, 28, 22, 87, 38], we summarize our contributions as follows:
@@
    We develop an RKHS regularized least-squares approach to learning general ODEs directly from sparse and partial observations. (see Section 2).
@@ 1.3 Literature Review
Learning continuous time dynamical systems from time series data has been a prominent problem in system identification [5, 57, 46]. Seminal works using symbolic regression [9, 82] through genetic programming sought to discover physical laws from observational data. The works [97, 12, 78, 81] developed and popularized the use of a sparsity prior over a feature library to promote parsimony in the discovered model, drawing on ideas from compressive sensing [24, 14, 15]. Bayesian approaches have also been considered, which allow for uncertainty quantification and motivate new strategies for model selection [102, 64, 101, 37, 65, 66].
@@
Many of these works consider a two-step framework which first estimates time-derivatives, then performs regression to estimate the dynamics. These approaches are straightforward to implement and often succeed when measurements of the process are sufficiently dense but fail when the data is scarcely sampled and noisy. These approaches first solve a smoothing problem to estimate the state and its derivatives, using methods such as total variation regularization [19], kernel regression [58], Savitzky–Golay filters, smoothing splines [96], or Kalman smoothing [86]. These approaches inevitably trade bias for robustness: as shown in [93], noise-resistant estimators lie on a Pareto frontier balancing error magnitude and error correlation due to implicitly assuming an inaccurate underlying dynamical model. An alternative to estimating derivatives, but that still lies within a two-step framework is to adopt a weak-form formulation [62], which avoids explicit differentiation by matching integrals against test functions. However, this still requires estimating integrals from noisy and scarcely sampled data, a problem that remains challenging despite being more well posed than differentiation. Consequently, any two-step approach to learning dynamics is inherently limited by the absence of a proper dynamical model in its initial estimation stage.
@@
To address limitations of two-step approaches, dynamics-informed approaches jointly estimate the dynamics and the state, allowing each to reinforce the other. These include Physics-Informed Neural Networks-Sparse Regression (PINN-SR) [22] and Physics-Informed Spline Learning (PiSL) [87] which use least squares collocation applied to neural networks and splines respectively, while obtaining sparsity in the coefficients of the ODE by alternating between applying variants of gradient descent and sequentially thresholded ridge regression. [40] takes a least squares collocation approach using RKHS methods, while also representing the unknown differential equation in an RKHS. The ODR-BINDy method in [28] applies a least squares relaxation to a finite difference discretization, and handles feature selection by greedily maximizing a Bayesian evidence over models defined by feature sets. [76] also applies a least squares relaxation to a finite difference discretization, but optimizes using a variational annealing approach, and focuses on identifying dynamics with hidden variables. Notably, our objective function (8) is closely related to that of [28, 76, 87, 51] and [22], though we apply different discretization, variable selection, and optimization strategies. [38] applies a finite difference discretization but imposes hard ODE constraints by solving the KKT system associated to the discretized problem. They promote sparsity via $\ell_{p}$ regularization, and optimize using iteratively reweighted least squares.
@@
Aside from [38], which considers decreasing sampling frequencies, and [76], which considers hidden variables, none of these works deal with very low sampling frequencies, partial observations through general linear measurements, and higher order dynamics in the context of ODEs. These issues are, however, considered heavily in [22] and [40] in the context of PDEs, where partial information on the state (being infinite dimensional) and higher order derivatives are fundamental to the problem. [38] crucially point out that decoupling the level of discretization in their finite difference stencil from the sampling rate of the data is required for mitigating bias in estimating the ODE when the sampling rate of the data is low, and demonstrate this in a restricted setting on the Van der Pol system. We generally consider sampling rates over an order of magnitude lower than other works and flexibly handle partial observations and higher order ODEs.
@@ 1.5 Notation
2 Joint state estimation and system identification
@@ 2.1 The abstract formulation
Our approach centers around the joint computation of the estimated state $\widehat{\boldsymbol{x}}(t)$, henceforth called state estimation, as well as the right hand side $\widehat{f}$, henceforth called system identification. We will require that the pair $\widehat{\boldsymbol{x}}$ and $\widehat{f}$ are self consistent, in the sense that $\widehat{\boldsymbol{x}}$ should (approximately) satisfy the ODE defined by $\widehat{f}$:
@@
To this end, we impose self-consistency using a least squares penalty method [95, 53, 3, 94, 100, 72], which is discretized via collocation in an RKHS. Imposing that $\widehat{f}$ has a sparse representation in the library $\Phi=\{\phi_{1},\ldots,\phi_{Q}\}$, we adopt a parametrized notation for how it operates on $\boldsymbol{x}$ and its derivatives:
@@
For instance, if we consider the first order constant coefficient linear ODE $\frac{d}{dt}\boldsymbol{x}(t)=\boldsymbol{A}\boldsymbol{x}$, then $D[\boldsymbol{x}](t)=\boldsymbol{x}$, and we would be looking for $\boldsymbol{\theta}=\boldsymbol{A}$.
@@
    Because we handle sparsity through the feature selection operator $\mathcal{S}$ we do not need to apply regularization to $\boldsymbol{\theta}$ to promote sparsity. In principle, however, the Frobenius norm $\frac{\mu}{2}\|\boldsymbol{\theta}\|_{F}^{2}$ could be replaced by a more general sparsity promoting regularizer $R(\boldsymbol{\theta})$ which would supplant $\mathcal{S}$.
@@
then the pair $(\boldsymbol{x}^{*},\boldsymbol{\theta}^{*})$ is a fixed point of (6). This is a strong condition, but is likely to hold whenever $\mathcal{S}$ is a reasonably good sparsifier that is consistent with the collocation loss $\mathcal{L}_{C}$. This is also a desirable property in itself; we want for our estimate fixed to the active set to be the best estimate possible on that active set.
The ADO algorithms of [87, 22] are exactly of the form (6) and motivate it as a surrogate to $\ell_{0}$ minimization. Similarly, the greedy evidence maximization in [28] is related in that it solves a sequence of fixed-support nonlinear least squares problems, though in each iteration, separate subproblems are solved for each candidate feature to remove. This has the benefit of minimizing a fixed objective function, and ensuring convergence due to the monotone feature removal process. This is discussed in detail algorithmically in Section 3.1 and Section 3.3.
@@
To represent the $d$-dimensional trajectory $\boldsymbol{x}:[0,T]\to\mathbb{R}^{d}$, we work in a vector-valued RKHS $\mathcal{X}$ of functions $0,T]\mapsto\mathbb{R}^{d}$ generated by the $d$-fold Cartesian product of an RKHS $\mathcal{H}_{k}$, associated to a symmetric positive definite kernel $k$. This is equivalent to modeling each coordinate $x_{i}(t)$ of the trajectory $\boldsymbol{x}(t)$ separately in the RKHS $\mathcal{H}_{k}$, or by taking $\mathcal{X}$ to be the vector-valued RKHS with matrix valued kernel $K(t,t^{\prime})=k(t,t^{\prime})I_{d}$. We will require that for derivative orders j=0,,pj=0,...,p, and times $t\in[0,T]$, the derivative evaluation map $\boldsymbol{x}\mapsto\frac{d^{j}}{dt^{j}}\boldsymbol{x}(t)$ is a bounded linear operator on $\mathcal{X}$ (see Section A.2). It suffices to take $k$ such that $\mathcal{H}_{k}$ is compactly embedded in $C^{p}([0,T])$, which is satisfied if $k\in C^{p,p}([0,T]\times[0,T])$ [104].
We choose the kernel $k$ to be the sum of a high-order Matérn kernel [74] and a constant kernel to account for systems where the time-average is nonzero. More precisely, with $h_{\nu}$ denoting the $\nu$-order Matérn covariance function and $\ell$ as a length scale parameter, we use kernels of the form
@@
where $K_{\nu}$ is the modified Bessel function of the second kind, and $c_{0},c_{1}>0$. These kernels induce RKHSs which satisfy the boundedness condition of the derivative evaluation maps whenever $\nu>p+\frac{1}{2}$. See Section A.2 for more details on kernels and smoothness.
As a preliminary step in our model, we choose the kernel hyperparameters $c_{0},c_{1},\ell$, and the variance of the observation noise $\sigma^{2}=\mathbb{V}\left[\epsilon\right]$ using maximum likelihood and treat the distribution of $\epsilon$ as Gaussian, as is common in Gaussian process regression; see for example [74]. From there, $\|\boldsymbol{x}\|_{\mathcal{X}}^{2}$ corresponds to the regularization induced by the RKHS norm associated to this fitted Gaussian process.